QUESTION 3: CAPITAL BUDGETING [30 MARKS]
Swee Rien Roofing Materials, Inc., is considering two mutually exclusive projects, each with an initial investment of RM1,500,000. The company’s board of directors has set a maximum 4-year payback requirement and has set its cost of capital at 9.50 percent. The cash inflows associated with the two projects are shown in the following table.

Cash inflows (CFt)

Year

Project A (RM)

Project B (RM)

1

450,000

750,000

2

450,000

600,000

3

550,000

300,000

4

400,000

350,000

5

450,000

250,000

6

450,000

300,000


a) Find the payback period for each project.
(6 marks)
b) Calculate the NPV of each project at 9.50 percent.
(7 marks)
c) Compute the Profitability Index.
(5 marks)
d) Derive the IRR of each project.
(4 marks)
e) Rank the projects by each of the techniques used. Make and justify a recommendation.
(8 marks)
Answer #1

A) payback period for Project A

Cash inflows (CFt) (in RM) (in RM)
Year cost Cash flow working cost yet to be recovered
0 1500000 0 0
1 4,50,000 (1500000-4500000) 10,50,000
2 4,50,000 (1050000-450000) 6,00,000
3 5,50,000 (600000-550000) 50,000
4 4,00,000 (50000-400000) -3,50,000
5 4,50,000
6 4,50,000
initial cost = 1500000
as we can see that the cost to be recovered turns negative in the 4th year
that means that the cost is recovered during 3rd and 4th year
Now, Payback period = years before full recovery + (Un recovered investment at start of the year/Cash flow during the year)
Payback period =3 +(50000/400000)
Payback period =3.125 years
Cash inflows (CFt)
Year cost Cash flow working cost yet to be recovered
0 1500000 0 0
1 7,50,000 (1500000-750000) 7,50,000
2 6,00,000 (750000-600000) 1,50,000
3 3,00,000 (150000-300000) -1,50,000
4 3,50,000
5 2,50,000
6 3,00,000
initial cost = 1500000
as we can see that the cost to be recovered turns negative in the 4th year
that means that the cost is recovered between 2nd and 3rd year
Now, Payback period = years before full recovery + (Unrecovered investment at start of the year/Cash flow during the year)
Payback period = 2+ (150000/300000)
Payback period =2.5 years

B).

Cash inflows (CFt) of Project A
R= 9.5%
Year COST Cash Flow working Discount factor = 1/(1+r)^n Discounted cash Flow= Cash flow * discount value
0 1500000 1/ (1+0.095)^0                        1.00                        -  
1 4,50,000 1/ (1+0.095)^1                        0.91        4,10,958.90
2 4,50,000 1/ (1+0.095)^2                        0.83        3,75,304.94
3 5,50,000 1/ (1+0.095^3                        0.76        4,18,909.62
4 4,00,000 1/ (1+0.095)^4                        0.70        2,78,229.72
5 4,50,000 1/ (1+0.095)^5                        0.64        2,85,852.45
6 4,50,000 1/ (1+0.095)^6                        0.58        2,61,052.47
Total      20,30,308.09
Net Present Value = Cash Inflows from Investments – Cost of Investments
NPV= 2030308.09-1500000
NPV= RM 530308.09
Cash inflows (CFt) of Project B
R= 9.5%
Year COST Cash Flow working Discount factor = 1/(1+r)^n Discounted cash Flow= Cash flow * discount value
0 1500000 1/ (1+0.095)^0                        1.00                        -  
1 7,50,000 1/ (1+0.095)^1                        0.91        6,84,931.51
2 6,00,000 1/ (1+0.095)^2                        0.83        5,00,406.58
3 3,00,000 1/ (1+0.095^3                        0.76        2,28,496.16
4 3,50,000 1/ (1+0.095)^4                        0.70        2,43,451.00
5 2,50,000 1/ (1+0.095)^5                        0.64        1,58,806.92
6 3,00,000 1/ (1+0.095)^6                        0.58        1,74,034.98
Total      19,90,127.14
Net Present Value = Cash Inflows from Investments – Cost of Investments
NPV= 1990127.14-1500000
NPV= RM 4,90,127.14

C).

Project A
Profitability index= Present value of future cash flow/ Initial investment
as seen in the above table,
Present value of future cash flow = RM 20,30,308.09
initial investment = RM 15,00,000
Profitability index= 2030308.09/1500000
PI= 1.3535
Project B
Profitability index= Present value of future cash flow/ Initial investment
as seen in the above table,
Present value of future cash flow = RM 19,90,127.14
initial investment = RM 15,00,000
Profitability index= 1990127.14/1500000
PI= 1.3267

D)

For calculating IRR, NPV= 0
So, we calculate the NPV at a assumed rate , IRR, which will provide us neither profit nor loss.
IRR is calculated through trial and error
For Project A
if we take IRR as 20%,
NPV= -1500000+ 450000/(1+0.2)^1+ 450000/(1.2)^2+ 550000/1.2^3+ 400000/1.2^4+ 450000/1.2^5+ 450000/1.2^6
= -1500000+1530237 = 30237
which is close to 1500000
now lets try 20.5%
NPV= -1500000+ 450000/(1+0.205)^1+ 450000/(1.205)^2+ 550000/1.205^3+ 400000/1.205^4+ 450000/1.205^5+ 450000/1.205^6
NPV= 11,531
similarly,Now we can try R as 20.8%
NPV= 487.81
So, we have estimated IRR= 20.8% , where NPv is closest to 0

For Project B

For calculating IRR, NPV= 0
So, we calculate the NPV at a assumed rate , IRR, which will provide us neither profit nor loss.
IRR is calculated through trial and error
For Project B
if we take IRR as 20%,
NPV= -1500000+ 750000/(1+0.2)^1+ 600000/(1.2)^2+ 300000/1.2^3+ 350000/1.2^4+ 250000/1.2^5+ 300000/1.2^6
= 1585005 -1500000= 85005
which is not too close to 1500000
now lets try 22%
NPV= -1500000+ 750000/(1+0.22)^1+ 600000/(1.22)^2+ 300000/1.22^3+ 350000/1.22^4+ 250000/1.22^5+ 300000/1.22^6
NPV= 24556
similarly,Now we can try R as 23%
NPV= -4090
So we will noe take a value close to 30%, lets take 22.85%
NPV= 142
So, we have estimated IRR= 22.8% , where NPV is closest to 0

E)

Project A Project B Rank
Payback period 3.125 years 2.5 years 1.Project B 2. Project A
NPV RM 530308 RM 490127 1.Project A 2. Project B
PI 1.35 1.32 1.Project A 2. Project B
IRR 20.80% 22.80% 1.Project B 2. Project A
The NPV method uses a reinvestment rate which is close to its cost of capital and the reinvestment assumptions of the NPV method much are more realistic than those with the IRR method. Also NPV is more realistic in non even cash flow senarios, because uneven cash flow leads to Many IRR and it is difficult to determine the exact rate,
Thus, NPV is a better indicator of evaluation than IRR
So We recommend Project A , as it has higher NPV and PI as compared to project B.